# Non-Asymptotic Analysis of Efficiency in Conformalized Regression

### v1

@@ Appendix B Proofs of Results in CQR
To prove Theorem 3.1, we first show that $\ell_{\gamma}\left(\theta\right)$ in (4) is strongly convex and smooth with respect to $\theta^{*}(\gamma)$, as stated below in Proposition B.1. The proof of Proposition B.1 further relies on Lemma B.2 and Lemma B.3 for the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$.
Proposition B.1.
@@
To prove this proposition, we first need Lemma B.2 and Lemma B.3 to calculate the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$. By Lemma B.2,
@@
By Lemma B.3, $\nabla^{2}\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]$. By Assumption 3.3, $\forall v\in\mathbb{R}^{d}$,
@@
By Assumption, $\mathbb{E}\left[\|X\|_{2}\right]\leq\sqrt{\mathbb{E}\left[\|X\|^{2}\right]}<\infty$. Then, by Lemma B.2,
@@
To prove the lemma, the key point is to show that the interchange of differentiation and expectation is valid, as in the proof of Lemma B.2.
@@
With Proposition B.1, we are ready to apply Theorem B.4 for SGD and get Corollary B.1.
Theorem B.4 (Section 3 in Rakhlin et al. (2012)).
@@
Corollary B.1 (Upper Bound of Extra Loss).
@@
In this proof, we denote $\theta_{n}\left(\gamma\right)$ by $\theta_{n}$ for simplicity. By Lemma B.2, $\nabla\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}\right)X\right]$. Then,
@@
The corollary then follows from Proposition B.1 and Theorem B.4. ∎
Now we are ready to prove Theorem 3.1. In this proof, we denote $\theta_{n}\left(\gamma\right),\theta^{*}\left(\gamma\right)$ by $\theta_{n},\theta^{*}$, respectively, for simplicity. By Proposition B.1,
@@
Hence, by Corollary B.1, $\mathbb{E}_{\theta_{n}}[\|\theta_{n}-\theta^{*}\|_{\Sigma}^{2}]\leq\frac{2}{f_{\min}}\mathbb{E}_{\theta_{n}}[\left(\ell\left(\theta_{n}\right)-\ell\left(\theta^{*}\right)\right)]\leq\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{3}f_{\min}^{2}n}$.
@@
B.2 Proof of Proposition B.5
Proposition B.5 (Population quantile of the score).
@@
The proof of Proposition B.5 relies on the following lemma.
Lemma B.6.
@@
By Lemma B.6, $\left|q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\right|\leq\Delta\left(X,\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)$. Then,
@@
This completes the proof of Proposition B.5.
B.3 Proof of Proposition B.7
Proposition B.7 (Population finite-sample score-quantile gap).
@@
To prove Proposition B.7, we first need the following critical proposition:
Proposition B.8.
@@
and by Proposition B.5, it also holds that
@@
The proof of Proposition B.7 also relies on the following useful lemma.
Lemma B.9.
@@
With Proposition B.8, we apply Lemma B.9 and get Lemma B.10.
Lemma B.10.
Under the same setting of Proposition B.8, if the event in Proposition B.8 occurs, and if $m>\frac{4}{f_{\min}\beta}$, then it holds that
@@
Since $\varepsilon_{n}<\beta/4$, from Proposition B.8, with probability at least $1-\delta$, for $s$ with $|s|<3\beta/4$, $f_{S\mid\vartheta_{n}}\left(s\mid\vartheta_{n}\right)\geq 2f_{\min}$, and $|q_{1-\alpha}|<\beta/4$. In this case, $r_{0}:=\min\{q_{1-\alpha}+3\beta/4,3\beta/4-q_{1-\alpha}\}>\beta/2$. If $m>\frac{4}{f_{\min}\beta}$, then $|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2}{m}<2f_{\min}\frac{\beta}{4}<2f_{\min}\frac{\beta}{2}<2f_{\min}\cdot r_{0}$. Then by Lemma B.9, $|q_{\left(1-\alpha\right)_{m}}|\leq 3\beta/4$, $f_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)\right)\geq 2f_{\min}$, and $|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\frac{|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|}{2f_{\min}}<\frac{1}{f_{\min}m}\leq\beta/4$. Hence, $|q_{\left(1-\alpha\right)_{m}}|\leq|q_{1-\alpha}|+|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\beta/4+\beta/4=\beta/2$. ∎
Notice that $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|$ is bounded by $2R$. Let $V$ denote the event in Proposition B.8, and $V^{c}$ its complement. Then, by Lemma B.10,
@@
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition B.7.
B.4 Proof of Proposition B.11
Proposition B.11 (Empirical score-quantile concentration).
@@
To prove Proposition B.11, we first prove the following lemma:
Lemma B.12.
Under the same setting of Lemma B.10, if the high probability event $V$ in Proposition B.8 occurs, for any $u\in[0,\beta/4]$, if
@@
For simplicity, in the proof we denote $q_{p}\left(S\mid\vartheta_{n}\right)$ by $q_{p}$. By Lemma B.10, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}+u|\leq 3\beta/4$. Hence, in this case,
@@
Lemma B.13 (Dvoretzky–Kiefer–Wolfowitz Inequality (Dvoretzky et al., 1956; Massart, 1990)).
@@
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma B.13),
@@
Thus, by Lemma B.12, given that the event $V$ occurs,
@@
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition B.11.
B.5 Proof of Theorem 3.2
@@
Appendix C Proofs of Results in CMR
@@
Thus, the expectation is decomposed into three parts as follows, and we will upper bound each of them in Proposition C.4, C.3, and C.1:
@@
C.1 Proof of Proposition C.1
Proposition C.1.
@@
C.2 Proof of Proposition C.2
Proposition C.2.
@@
C.3 Proof of Proposition C.3
Proposition C.3.
@@
According to Lemma B.9, since $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}<\frac{\beta}{4}$ by Proposition C.2, the distance from $\mathcal{I}^{c}$ is $r_{0}>\frac{\beta}{2}$. Thus, by Lemma B.9, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\leq\frac{1}{f_{\min}m}<\frac{\beta}{4}$, and hence, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-\zeta|<\frac{\beta}{2}$.
@@
C.4 Proof of Proposition C.4
Proposition C.4.
@@
The proof of Proposition C.4 is essentially the same as the proof of Proposition B.11. We include here for completeness.
@@
Lemma C.5.
In CMR, under the same setting of Proposition C.2, if the high probability event $V$ in Proposition C.2 occurs, for any $u\in[0,\beta/4]$, if
@@
For simplicity, in the proof we denote $q_{p}\left(S\mid\check{\theta}_{n}\right)$ by $q_{p}$. By Proposition C.3, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-\zeta-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}-\zeta+u|\leq 3\beta/4$, i.e., $q_{\left(1-\alpha\right)_{m}}-u\in\mathcal{I}$ and $q_{\left(1-\alpha\right)_{m}}+u\in\mathcal{I}$ for $\mathcal{I}$ defined in Proposition C.2. Hence, in this case,
@@
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma B.13),
@@
Thus, by Lemma B.12, given that the event $V$ occurs,
@@
Picking $\delta=\frac{257\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$ completes the proof of Proposition C.4. ∎
Appendix D Additional Experiments on Synthetic Data
D.1 Data Generation in Section 6
@@
D.2 Validating Regime of $\mathcal{O}(1/(n\alpha^{2}))$
@@
Appendix E Experiments on Real-World Data

### v2

@@ Appendix A Proofs of Results in CQR
To prove Theorem 3.1, we first show that $\ell_{\gamma}\left(\theta\right)$ in (4) is strongly convex and smooth with respect to $\theta^{*}(\gamma)$, as stated below in Proposition A.1. The proof of Proposition A.1 further relies on Lemma A.2 and Lemma A.3 for the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$.
Proposition A.1.
@@
where $\|\cdot\|_{\Sigma}\$denotes the $\Sigma$-induced norm, i.e., $\|\theta\|_{\Sigma}:=\sqrt{\theta^{\top}\Sigma\theta}$.
@@
To prove this proposition, we first need Lemma A.2 and Lemma A.3 to calculate the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$. By Lemma A.2,
@@
By Lemma A.3, $\nabla^{2}\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]$. By Assumption 3.3, $\forall v\in\mathbb{R}^{d}$,
@@
By Assumption, $\mathbb{E}\left[\|X\|_{2}\right]\leq\sqrt{\mathbb{E}\left[\|X\|^{2}\right]}<\infty$. Then, by Lemma A.2,
@@
To prove the lemma, the key point is to show that the interchange of differentiation and expectation is valid, as in the proof of Lemma A.2.
@@
With Proposition A.1, we are ready to apply Theorem A.4 for SGD and get Corollary A.1.
Theorem A.4 (Section 3 in rakhlin2012making).
@@
Corollary A.1 (Upper Bound of Extra Loss).
@@
In this proof, we denote $\theta_{n}\left(\gamma\right)$ by $\theta_{n}$ for simplicity. By Lemma A.2, $\nabla\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}\right)X\right]$. Then,
@@
The corollary then follows from Proposition A.1 and Theorem A.4. ∎
Now we are ready to prove Theorem 3.1. In this proof, we denote $\theta_{n}\left(\gamma\right),\theta^{*}\left(\gamma\right)$ by $\theta_{n},\theta^{*}$, respectively, for simplicity. By Proposition A.1,
@@
Hence, by Corollary A.1, $\mathbb{E}_{\theta_{n}}[\|\theta_{n}-\theta^{*}\|_{\Sigma}^{2}]\leq\frac{2}{f_{\min}}\mathbb{E}_{\theta_{n}}[\left(\ell\left(\theta_{n}\right)-\ell\left(\theta^{*}\right)\right)]\leq\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{3}f_{\min}^{2}n}$.
@@
A.2 Proof of Proposition A.5
Proposition A.5 (Population quantile of the score).
@@
The proof of Proposition A.5 relies on the following lemma.
Lemma A.6.
@@
By Lemma A.6, $\left|q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\right|\leq\Delta\left(X,\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)$. Then,
@@
This completes the proof of Proposition A.5.
A.3 Proof of Proposition A.7
Proposition A.7 (Population finite-sample score-quantile gap).
@@
To prove Proposition A.7, we first need the following critical proposition:
Proposition A.8.
@@
and by Proposition A.5, it also holds that
@@
The proof of Proposition A.7 also relies on the following useful lemma, which is a classical result (bobkov2019one). We include the proof here for completeness.
Lemma A.9.
@@
With Proposition A.8, we apply Lemma A.9 and get Lemma A.10.
Lemma A.10.
Under the same setting of Proposition A.8, if the event in Proposition A.8 occurs, and if $m>\frac{4}{f_{\min}\beta}$, then it holds that
@@
Since $\varepsilon_{n}<\beta/4$, from Proposition A.8, with probability at least $1-\delta$, for $s$ with $|s|<3\beta/4$, $f_{S\mid\vartheta_{n}}\left(s\mid\vartheta_{n}\right)\geq 2f_{\min}$, and $|q_{1-\alpha}|<\beta/4$. In this case, $r_{0}:=\min\{q_{1-\alpha}+3\beta/4,3\beta/4-q_{1-\alpha}\}>\beta/2$. If $m>\frac{4}{f_{\min}\beta}$, then $|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2}{m}<2f_{\min}\frac{\beta}{4}<2f_{\min}\frac{\beta}{2}<2f_{\min}\cdot r_{0}$. Then by Lemma A.9, $|q_{\left(1-\alpha\right)_{m}}|\leq 3\beta/4$, $f_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)\right)\geq 2f_{\min}$, and $|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\frac{|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|}{2f_{\min}}<\frac{1}{f_{\min}m}\leq\beta/4$. Hence, $|q_{\left(1-\alpha\right)_{m}}|\leq|q_{1-\alpha}|+|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\beta/4+\beta/4=\beta/2$. ∎
Notice that $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|$ is bounded by $2R$. Let $V$ denote the event in Proposition A.8, and $V^{c}$ its complement. Then, by Lemma A.10,
@@
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition A.7.
A.4 Proof of Proposition A.11
Proposition A.11 (Empirical score-quantile concentration).
@@
To prove Proposition A.11, we first prove the following lemma:
Lemma A.12.
Under the same setting of Lemma A.10, if the high probability event $V$ in Proposition A.8 occurs, for any $u\in[0,\beta/4]$, if
@@
For simplicity, in the proof we denote $q_{p}\left(S\mid\vartheta_{n}\right)$ by $q_{p}$. By Lemma A.10, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}+u|\leq 3\beta/4$. Hence, in this case,
@@
Lemma A.13 (Dvoretzky–Kiefer–Wolfowitz Inequality (dvoretzky1956asymptotic; massart1990tight)).
@@
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma A.13),
@@
Thus, by Lemma A.12, given that the event $V$ occurs,
@@
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition A.11.
A.5 Proof of Theorem 3.2
@@
Appendix B Proofs of Results in CMR
@@
Thus, the expectation is decomposed into three parts as follows, and we will upper bound each of them in Proposition B.4, B.3, and B.1:
@@
B.1 Proof of Proposition B.1
Proposition B.1.
@@
B.2 Proof of Proposition B.2
Proposition B.2.
@@
B.3 Proof of Proposition B.3
Proposition B.3.
@@
According to Lemma A.9, since $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}<\frac{\beta}{4}$ by Proposition B.2, the distance from $\mathcal{I}^{c}$ is $r_{0}>\frac{\beta}{2}$. Thus, by Lemma A.9, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\leq\frac{1}{f_{\min}m}<\frac{\beta}{4}$, and hence, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-\zeta|<\frac{\beta}{2}$.
@@
B.4 Proof of Proposition B.4
Proposition B.4.
@@
The proof of Proposition B.4 is essentially the same as the proof of Proposition A.11. We include here for completeness.
@@
Lemma B.5.
In CMR, under the same setting of Proposition B.2, if the high probability event $V$ in Proposition B.2 occurs, for any $u\in[0,\beta/4]$, if
@@
For simplicity, in the proof we denote $q_{p}\left(S\mid\check{\theta}_{n}\right)$ by $q_{p}$. By Proposition B.3, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-\zeta-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}-\zeta+u|\leq 3\beta/4$, i.e., $q_{\left(1-\alpha\right)_{m}}-u\in\mathcal{I}$ and $q_{\left(1-\alpha\right)_{m}}+u\in\mathcal{I}$ for $\mathcal{I}$ defined in Proposition B.2. Hence, in this case,
@@
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma A.13),
@@
Thus, by Lemma A.12, given that the event $V$ occurs,
@@
Picking $\delta=\frac{257\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$ completes the proof of Proposition B.4. ∎
Appendix C Additional Experiments on Synthetic Data
C.1 Data Generation in Section 6
@@
C.2 Validating Regime of $\mathcal{O}(1/(n\alpha^{2}))$
@@
C.3 Training via Alternative Optimizers
To demonstrate that our analytical framework extends directly to alternative optimization algorithms by substituting the corresponding estimation error rate, Figure 6 reports the empirical results obtained using SGD with heavy-ball momentum (polyak1964some). Theoretically, SGD with momentum achieves the same convergence rate as vanilla SGD, up to improved constants. According to Remark 3.3, the efficiency with SGD with momentum scales in the same order as SGD. Consistent with this prediction, the empirical results show that the phase transition phenomenon identified in our analysis persists under SGD with momentum as well. Specifically, in Figure 6 (c), the slope of curves in Figure 6 (a) changes from 1-1 to 0.5-0.5 as $\alpha$ increases.
Moreover, to demonstrate that our theoretical insights are not tied to optimizers with established convergence guarantee, Figure 7 reports the empirical results obtained using AdamW (loshchilov2019decoupled). From Figure 7 (c), we observe the phase transition phenomenon, where the slope of curves in Figure 7 (a) changes from 1-1 to 0.5-0.5 as $\alpha$ increases.
Figure 6: The length deviation of conformalized quantile regression with training via SGD with momentum.
Figure 7: The length deviation of conformalized quantile regression with training via AdamW.
C.4 Non-Linear Ground-Truth Quantile Functions
To empirically show that our theoretical insights extend beyond linear models, we conducted experiments in a setting where the ground-truth quantile functions are no longer linear. This is achieved by applying Gaussian convolution kernels to the original linear conditional probability density functions, thereby introducing controlled non-linear distortion. As shown in Figure 8, even with this non-linear distortion, the phase transition phenomenon persists, indicating that our theoretical insights remain valid in a broader setting.
Figure 8: The length deviation of conformalized quantile regression on the data distribution with non-linear conditional quantile functions, where Gaussian convolution kernels are applied to the linear conditional probability density functions. We set $\sigma$ to be $0.1\times$ conditional quantile.
C.5 Alternative Loss Functions
To provide empirical evidence that similar efficiency scaling behavior persists for other convex models satisfying our assumptions, we report results using $\ell_{1}$-regularization during training in Figure 9 and Huber penalty (huber1964robust) during training in Figure 10. In both cases, the phase transition phenomenon remains clearly visible (the slope of (a) changes from 1-1 to 0.5-0.5 as $\alpha$ increases), further validating the generality of our theoretical insights.
Figure 9: The length deviation of conformalized quantile regression, training with $\ell_{1}$ regularization. We set the coefficient of the regularization term to be $0.001$.
Figure 10: The length deviation of conformalized quantile regression, training with Huber penalty. We set Huber $\delta$ to be $0.5$, and Huber $\lambda$ to be $0.1$.
Appendix D Experiments on Real-World Data
D.1 Statistics of Datasets
We list the statistics of multiple popular real world regression datasets used in this paper in Table 1. The empirical investigation of their distribution patterns are provided in Appendix D.4.